OpenStudy (anonymous):

can I factor a^2 - a7 - 15 as a(a-7)-15?

Mehek (mehek14):

-1 • (a7 - a2 + 15)

Mehek (mehek14):

you can factor it like that

OpenStudy (anonymous):

is a(a-7)-15 still acceptable?

Mehek (mehek14):

a(a-7)-15 = \(a^2-a7-15\) also, the remaining numbers, like 15 have to be in parenthesis

OpenStudy (anonymous):

Oh sorry, the question was suppose to be a^2-7a-15

OpenStudy (anonymous):

I edited it

Mehek (mehek14):

no that can't be factored

OpenStudy (anonymous):

wouldn't that equal a(a-7)-15 = a^2−a7−15?

Mehek (mehek14):

no

Mehek (mehek14):

because the variable outside the parenthesis has to be common in each term -15 does not contain a

OpenStudy (anonymous):

I see...thank you for you assistance

OpenStudy (anonymous):

your*

Mehek (mehek14):

yw =]